8. (a) Let X X Xn

, , , 1 2 K be independently and identically distributed b ,1( p) random 

variables. Obtain a confidence internal for p using Chebychev’s inequality. 


1
Expert's answer
2021-03-26T17:11:51-0400

For sample mean

"\overline{X}=\frac{1}{N}\sum x_k"

using Chebishev inequality:

"P(|\overline{X}-\mu|\geq b)\leq\frac{\sigma^2}{Nb^2}"